OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. -x - 3y = 9.

OpenStudy (anonymous):

For x intercept put y = 0 And for y intercept put x = 0..

hartnn (hartnn):

Get x-intercept by putting y=0 in that equation Get y-intercept by putting x=0 in that equation can u do that ?

OpenStudy (anonymous):

Ok, I then have to graph the equation, so will it be at 0 for both y and x?

hartnn (hartnn):

no,and no need to graph....For x-intercept,u put y=0 like this: -x-3(0)=9 x=-9 is the x-intercept can u do the same for y-intercept??

OpenStudy (anonymous):

Firstly find x and y intercepts..

OpenStudy (anonymous):

Ok, I do have to submit a graph in the dropbox as part of my quiz.

OpenStudy (anonymous):

I'm so confused.

OpenStudy (anonymous):

Firstly find the y intercept.. x intercept has been found by hartnn..

OpenStudy (anonymous):

After that you have to plot the points on graph..

OpenStudy (anonymous):

Ok well I guess I will figure it out eventually, thanks.

OpenStudy (anonymous):

Is it -0 - 3y =9?

OpenStudy (anonymous):

Anyone there?

hartnn (hartnn):

what will -0 - 3y =9 give you the value of y? that will be y-intercept..

OpenStudy (anonymous):

y=0?

OpenStudy (anonymous):

I'm sorry if I seen stupid but this is all very new to me.

OpenStudy (anonymous):

do u know how to solve linear equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I don't know how to do any of this.

OpenStudy (anonymous):

hmm...then u need to learn that before solving this question. try this link. see if it helps. http://www.studygs.net/equations.htm if not, let us know. we'll try and help u out.

hartnn (hartnn):

-0 - 3y =9 - 3y =9 \[\frac {-3y}{-3}=\frac{9}{-3}\] y=-3 does this make sense??

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so now u have x-intercept (-9,0) y- intercept (0,-3) ok ...? now can u plot such points in the graph??

OpenStudy (anonymous):

Yes I can.....Thanks for all of you alls help and I apologize for giving you a headache

hartnn (hartnn):

no problem,but could you graph that equation?

OpenStudy (anonymous):

Yes that much I can do lol